# Federated Cubic Regularized Newton Learning with Sparsification-amplified Differential Privacy

### v1

@@ Abstract
This paper investigates the use of the cubic-regularized Newton method within a federated learning framework while addressing two major concerns that commonly arise in federated learning: privacy leakage and communication bottleneck. We introduce a federated learning algorithm called Differentially Private Federated Cubic Regularized Newton (DP-FCRN). By leveraging second-order techniques, our algorithm achieves lower iteration complexity compared to first-order methods. We also incorporate noise perturbation during local computations to ensure privacy. Furthermore, we employ sparsification in uplink transmission, which not only reduces the communication costs but also amplifies the privacy guarantee. Specifically, this approach reduces the necessary noise intensity without compromising privacy protection. We analyze the convergence properties of our algorithm and establish the privacy guarantee. Finally, we validate the effectiveness of the proposed algorithm through experiments on a benchmark dataset.
@@ 1 Introduction
As big data continues to grow and awareness of privacy issues increases, conventional centralized methods for optimizing model parameters encounter substantial challenges. To deal with these challenges, federated learning (FL) has become a promising approach. FL allows multiple computing devices to collaborate in a distributed paradigm, under the coordination of the central server, without the need to share their local data, to optimize a shared large model. FL has found wide applications such as robotics [1], finance [2], and autonomous driving [3].
@@
The prevailing algorithm used in FL is based on stochastic gradient descent (SGD), usually called Fed-SGD [4]. In this approach, each client trains the local model on its dataset using SGD and then uploads the gradient to the central server. The server then averages all the local gradients and performs a gradient step. Although first-order methods like SGD are the currently preferred choice for FL, they usually suffer from slow convergence, which can impede the deployment of systems that require fast and efficient processing, such as autonomous vehicles where timely and accurate predictions are vital. Newton’s technique, a second-order method, is renowned for its fast convergence in optimization. However, incorporating second-order methods into FL is highly challenging. The key obstacle lies in the non-linear nature of aggregating the solutions from local optimization problems for second-order approximation, unlike the straightforward gradient aggregation. This complexity is evident in recent algorithms like GIANT [5].
@@
While aggregating the local Hessians is possible in theory, uploading the Hessian matrices at each round incurs significant communication costs. Moreover, even without matrix transmission, communication efficiency remains a critical bottleneck in FL. For instance, if clients are some mobile devices, they usually have limited communication bandwidth. For example, the language model GPT-3 [6] has billions of parameters, and thus, it is impractical to transmit them directly. Traditional first-order optimization methods employ various techniques to improve communication efficiency, such as communication compression [7], event-based transmission [8, 9], and partial participation [10]. As for faster second-order approaches, Safaryan et al. [11] have recently introduced a family of federated Newton learning methods that facilitate general contractive compression operators for matrices and partial participation, thereby reducing communication costs. Liu et al. [12] developed a distributed Newton’s method with improved communication efficiency and achieved the super-linear convergence result. Fabbro et al. [13] presented a Newton-type algorithm to accelerate FL and considered communication constraints.
@@
In addition to slow convergence and communication costs in traditional FL, privacy leakage is another significant concern. That the data is stored locally on the client does not alone offer adequate privacy protection. For instance, recent inference attacks in [14, 15, 16] have demonstrated that sharing local model updates or gradients between clients and the server can result in privacy breaches. Sharing second-order information can also pose privacy concerns as it often encompasses the client’s data. For example, Yin et al. [17] have shown that even the eigenvalues of the Hessian matrix can disclose some critical information from original input images. Thus, it is imperative to preserve privacy in FL. Differential privacy (DP), introduced by Dwork [18], has been the de facto standard among various privacy-preserving frameworks, primarily because of its effectiveness in data analysis tasks and strong privacy guarantee. In differentially private Fed-SGD, the gradient is typically augmented with Gaussian noise to achieve DP [19, 20, 21, 22]. It is shown that due to the composition of DP, the required noise level at each iteration is always influenced by the number of iterations. Recently, Ganesh et al. [23] devised a differentially private optimization approach using second-order techniques, attaining $(\varepsilon,\delta)$-DP and the utility loss $O(d/\varepsilon^{2})$ for $d$ dimensional model. This utility loss bound is optimal, i.e., the best achievable for differentially private optimization [24, 25], and their algorithm achieves faster convergence. However, this method is restricted to the centralized setting. In FL, ensuring DP for second-order optimization is a novel area of exploration. We should address challenges such as integrating noise perturbation and reducing communication concurrently in federated Newton learning for each client, and enhancing the overall balance among privacy, accuracy, and communication efficiency.
@@
Motivated by the above observations, we aim to investigate federated Newton learning while jointly considering DP and communication issues in the algorithm design. Prior research predominantly considers DP and communication efficiency as separate entities [26, 27]. While some research has explored the joint trade-off among privacy, accuracy, and communication [28, 29], they tackled the communication and privacy in a cascaded fashion, i.e., their communication schemes do not directly impact privacy preservation. In contrast, our study investigates the interplay between communication and privacy guarantees. Although some recent studies have employed compression in uplink transmission to improve privacy [30, 31], these approaches are limited to first-order learning with slow convergence. Besides, Chen et al. [31] exclusively addressed central DP, which is less robust compared to privacy mechanisms at the client level. Specifically, we propose that each local machine runs a cubic regularized Newton method to update its model, with the addition of noise perturbation during the local computation process. In FL, since the local model update is usually sparse, we propose integrating perturbation with random sparsification to boost privacy preservation. Since sparsification remains only a subset of coordinates and sets values in other coordinates to zero, it reduces the sensitivity of updates to raw data, and results in lower privacy loss during each communication round. We show that the necessary noise intensity for DP is directly proportional to the number of transmitted coordinates. This implies that improved communication efficiency can reduce noise perturbation without compromising privacy. Furthermore, we illustrate that the complexity of our algorithm exhibits an exponential improvement compared to that of first-order methods. As the required noise intensity often depends on the number of iterations, this acceleration further reduces the noise intensity and enhances the balance between privacy and convergence trade-off.
The main contributions of our work are summarized as follows:
@@
    We develop a differentially private federated cubic regularized Newton learning algorithm (Algorithm 1), called DP-FCRN. By using second-order Newton methods, the proposed algorithm achieves rapid convergence. We exploit noise perturbation in local computations to guarantee privacy preservation (Algorithm 2). Moreover, we use sparsification to improve communication efficiency. Unlike previous studies that treat DP and communication efficiency as separate goals and neglect the impact of efficient communication on privacy [26, 27, 28, 29], we use the inherent characteristic of sparsification to enhance privacy.
@@
    We demonstrate the impact of sparsification on the balance between utility and privacy. Specifically, we show that the required noise intensity is reduced by a sparsification ratio (Theorem 1), which means that the proposed algorithm can exploit sparsified transmission to reduce the magnitude of Gaussian noise. Moreover, we conduct a non-asymptotic analysis and obtain the utility loss and the corresponding complexity (Theorem 2). The utility loss is optimal and the iteration complexity for achieving the optimal utility loss is improved over that of first-order methods.
@@
    We empirically evaluate our scheme on the benchmark dataset. The experiment results illustrate that our algorithm improves the model accuracy, and at the same time saves communication costs compared to Fed-SGD under the same DP guarantee.
@@ 2.1 Basic Setup
We make the following assumptions on the loss function of the training model:
@@
$l(\cdot,\zeta)$ is $L_{0}$-Lipschitz, $L_{1}$-smooth, and has $L_{2}$-Lipschitz Hessian for any $\zeta$.
Assumption 2.
$l(\cdot,\zeta)$ is $\mu$-strongly convex for any $\zeta$.
Assumption 3.
$\mathcal{X}$ has a finite diameter $D$.
From Assumptions 1 and 2, we infer that also $f_{i}(\cdot)$ and $f(\cdot)$ are $\mu$-strongly convex, $L_{0}$-Lipschitz, $L_{1}$-smooth, and have $L_{2}$-Lipschitz Hessian.
@@ 2.2 Newton Methods with Cubic Regularization
Newton methods [32] iteratively minimize a quadratic approximation of the function $f(\cdot)$ as
@@
The cubic regularized Newton method, initially introduced by Nesterov and Polyak [33], incorporates a second-order Taylor expansion with a cubic regularization term. In particular, the update is
@@
where $L_{2}$ is the Lipschitz Hessian constant in Assumption 1. The cubic upper bound of $f(x_{t})$ in (3) serves as a universal upper bound regardless of the specific characteristics of the objective function. However, the function to minimize in each step of (3) does not have a closed-form solution and it is limited to a centralized single node setting, which our algorithm addresses in a federated setting as discussed in Section 3.
@@ 2.3 Threat Model and DP
Local datasets typically contain sensitive user information. If problem (1) is addressed in an insecure environment, the leakage of information could pose a threat to personal and property privacy. This paper considers the following adversary model [34]:
@@
    the honest-but-curious central server that follows the given protocol but may be curious about clients’ private information and capable of inferring it according to the received messages.
@@
    certain clients colluding among themselves or with the central server to deduce private information about other legal clients.
@@
    an outside eavesdropper who can intercept all transmitted messages in the communication and training protocol but not actively destroy message transmissions.
Our considered adversary model is much stronger than some works that require a trusted third party [35, 36].
DP is a privacy concept widely adopted for quantifying privacy risk. It is a characteristic of a randomized algorithm $\mathcal{A}$ where the presence or absence of an individual in a dataset cannot be determined from the output of $\mathcal{A}$ [18]. Here, we present the formal definition of DP within the context of FL.
@@
Definition 2 states that the output distributions of neighboring datasets exhibit small variation. The factor $\varepsilon$ in (4) represents the upper bound of privacy loss by algorithm $\mathcal{A}$, and $\delta$ denotes the probability of breaking this bound. Therefore, a smaller $\varepsilon$ corresponds to a stronger privacy guarantee. The Gaussian mechanism is one of the commonly employed techniques to achieve DP.
@@
(Gaussian Mechanism [37]) A Gaussian mechanism $\mathcal{G}$ for a vector-valued computation $r:\zeta\to\mathbb{R}^{d}$ is obtained by computing the function $r$ on the input data $\zeta_{i}\in\zeta$ and then adding random Gaussian noise perturbation $\nu\sim\mathcal{N}(0,\sigma^{2}I_{d})$ to the output, i.e,
@@ 3 Main Algorithm
In general, there are two approaches for integrating sparsification and privacy in FL: (1) perturb first, then sparsify, and (2) sparsify first, then perturb. The first approach is direct and adaptable since sparsification preserves DP and integrates smoothly with all current privacy mechanisms. However, in the second approach, perturbation may compromise the communication savings achieved through sparsification. Furthermore, empirical observations suggest that the first approach outperforms the second one in some scenarios [38]. Therefore, we adopt the first approach in this study.
@@
Following the update of the local model parameter, each client uploads its model update $x_{i,t+1}-x_{t}$ to the server. To address the communication challenges in uplink transmissions, the random-$k$ sparsifier is employed to reduce the size of the transmitted message by a factor of $k/d$ [39]:
@@
As pointed out by Lacoste-Julien et al. [40], the output of Algorithm 2 can be computed online. Specifically, setting $z_{0}=\theta_{0}$, and recursively defining $z_{s}=\rho_{s}\theta_{s}+(1-\rho_{s})z_{s-1}$ for $s\geq 1$, with $\rho_{s}=\frac{2}{s+1}$. It is a straightforward calculation to check that $z_{\tau}=\sum_{s=0}^{\tau-1}\frac{2s}{\tau(\tau+1)}\theta_{s}$.
@@
Directly solving (3) requires substantial computational resources. Therefore, we leverage parallel cooperative solving by multiple clients to enhance learning efficiency. To address the absence of a closed-form solution in (5), we propose a local training approach for its resolution. Privacy preservation is ensured through noise perturbation during local training, while sparsification in uplink transmission not only reduces communication costs but also enhances privacy protection.

### v2

@@ Abstract
This paper explores the cubic-regularized Newton method within a federated learning framework while addressing two major concerns: privacy leakage and communication bottlenecks. We propose the Differentially Private Federated Cubic Regularized Newton (DP-FCRN) algorithm, which leverages second-order techniques to achieve lower iteration complexity than first-order methods. We incorporate noise perturbation during local computations to ensure privacy. Furthermore, we employ sparsification in uplink transmission, which not only reduces the communication costs but also amplifies the privacy guarantee. Specifically, this approach reduces the necessary noise intensity without compromising privacy protection. We analyze the convergence properties of our algorithm and establish the privacy guarantee. Finally, we validate the effectiveness of the proposed algorithm through experiments on a benchmark dataset.
@@
thanks: The work by W. Huo and L. Shi is supported by the Hong Kong RGC General Research Fund 16203723.
@@ 1 Introduction
As big data grows and privacy concerns rise, conventional centralized methods for optimizing model parameters encounter substantial challenges. Federated learning (FL) has emerged as a promising approach, allowing multiple devices to collaboratively optimize a shared model under the coordination of the central server, without sharing local data. FL has found applications in fields such as robotics (Yuan et al. (2024)) and autonomous driving (Nguyen et al. (2022)). The prevailing FL algorithm is Fed-SGD, based on stochastic gradient descent (SGD) (Lowy & Razaviyayn (2023)). In this approach, each client trains a local model using SGD and uploads the gradient to the central server, which averages the gradients and updates the model. However, such first-order methods suffer from slow convergence, which can hinder applications requiring fast processing, such as autonomous vehicles, where timely and accurate predictions are critical. Newton’s technique, a second-order method, offers faster convergence, but its integration into FL represents challenges. The key obstacle is the non-linear nature of aggregating solutions from local optimization problems for second-order approximations, in contrast to the simpler gradient aggregation used in first-order methods. This complexity is evident in recent algorithms like GIANT (Maritan et al. (2023)).
@@
While aggregating local Hessians is theoretically feasible, uploading Hessian matrices at each round incurs significant communication costs. Even without transmitting matrices, communication efficiency remains a critical bottleneck in FL. For instance, mobile devices, which are commonly used as clients, often have limited communication bandwidth. Traditional first-order optimization methods improve communication efficiency through techniques such as compression (Richtárik et al. (2021)) and event trigger (Huo et al. (2024)). Recent second-order methods, like federated Newton learning (Safaryan et al. (2021)), incorporate contractive compression and partial participation to reduce communication costs. Building on this, Zhang et al. (2024) further reduced communication rounds with lazy aggregation and enhanced convergence using cubic and gradient-regularized Newton methods. Liu et al. (2023) developed a distributed Newton’s method with improved communication efficiency and achieved super-linear convergence. Dal Fabbro et al. (2024) presented a Newton-type algorithm to accelerate FL while addressing communication constraints.
@@
In addition to slow convergence and communication costs, privacy leakage is another significant concern in FL. Simply storing data locally on clients does not guarantee adequate privacy. Recent inference attacks (Zhu et al. (2019); Liu et al. (2022)) show that sharing local model updates or gradients between clients and the server can result in privacy breaches. Sharing second-order information can also expose sensitive client data. For instance, Yin et al. (2014) showed that eigenvalues of the Hessian matrix can leak critical information from input images. In algorithms transmitting compressed Hessians (Safaryan et al. (2021); Zhang et al. (2024)), if compression does not significantly alter the eigenvalues, sensitive data may still be exposed. Therefore, privacy preservation is vital in second-order FL. Differential privacy (DP), introduced by Dwork (2006), has been the standard framework for privacy preservation due to its effectiveness in data analysis tasks. For example, Wang et al. (2023) explored the relationship between differential initial-value privacy and observability in linear dynamical systems. Wang et al. (2024) proposed a distributed shuffling mechanism based on the Paillier cryptosystem to enhance the accuracy-privacy trade-off in DP-preserving average consensus algorithms. In differentially private Fed-SGD, the gradient is typically augmented with Gaussian noise to achieve DP (Lowy & Razaviyayn (2023)). Due to the composition of DP, the required noise level is influenced by the number of iterations. Recently, Ganesh et al. (2024) devised a second-order, differentially private optimization method that achieves $(\varepsilon,\delta)$-DP with utility loss $O(d/\varepsilon^{2})$ for $d$-dimensional model, which is optimal, i.e., the best achievable, for differentially private optimization (Bassily et al. (2014); Kairouz et al. (2021)). However, this method is restricted to centralized settings. Ensuring DP for second-order optimization in FL remains a challenge, requiring the integration of noise perturbation with communication-efficient methods and addressing trade-offs between privacy, accuracy, and communication across clients.
@@
Motivated by the above observations, we aim to investigate federated Newton learning while jointly considering DP and communication issues in the algorithm design. Prior research predominantly considers DP and communication efficiency as separate entities (Li et al. (2022); Zhang et al. (2020)). While some research has explored the joint trade-off among privacy, accuracy, and communication (Mohammadi et al. (2021); Chen et al. (2021)), they tackled the communication and privacy in a cascaded fashion, i.e., their communication schemes do not directly impact privacy preservation. In contrast, our study investigates the interplay between communication and privacy guarantees. Although some recent studies have employed compression in uplink transmission to improve the trade-off (Hu et al. (2023); Chen et al. (2024)), these approaches are limited to first-order learning with slow convergence. Besides, Chen et al. (2024) exclusively addressed central DP, which is less robust compared to privacy mechanisms at the client level. Specifically, we propose that each local machine uses a cubic regularized Newton method for model updates, incorporating noise perturbation during local computation. In FL, where local model updates are typically sparse, we combine perturbation with random sparsification to enhance privacy. Sparsification reduces the sensitivity of updates to raw data by zeroing out some coordinates, thereby lowering privacy loss during communication. We show that the noise intensity required for DP is proportional to the number of transmitted coordinates, meaning improved communication efficiency can reduce the noise without compromising privacy. Furthermore, we illustrate that our algorithm’s iteration complexity exhibits an exponential improvement compared to first-order methods, further reducing noise intensity and enhancing the trade-off between privacy and convergence. Comparison of some related works with ours is shown in Table 1.
Table 1: Comparison of some existing works with ours
Work Efficient communication Optimization DP Impact of efficient communication on DP
Bassily et al. (2014); Kairouz et al. (2021) $\times$ First-order Client-DP $\times$
Chen et al. (2021, 2024) Compressed vectors First-order Central DP Increased compression did not improve privacy and reduced accuracy
Safaryan et al. (2021) Compressed matrices Second-order $\times$ $\times$
Zhang et al. (2024) Compressed matrices Second-order $\times$ $\times$
Ours Compressed vectors Second-order Client-DP Increased compression enhances privacy and improves accuracy
Our main contributions are summarized as follows:
@@
    We develop the DP-FCRN algorithm (Algorithm 1), which leverages second-order Newton methods for faster convergence. We exploit noise perturbation in local computations to guarantee privacy preservation (Algorithm 2) and use sparsification to improve communication efficiency. Unlike previous studies that treat DP and communication burden separately (Li et al. (2022); Zhang et al. (2020); Mohammadi et al. (2021); Chen et al. (2021)), we use the inherent characteristic of sparsification to simultaneously enhance both privacy and communication efficiency.
@@
    We analyze the impact of sparsification on the privacy-accuracy trade-off. Specifically, we show that sparsification reduces the required noise intensity (Theorem 1), allowing for lower Gaussian noise while maintaining privacy. We also conduct a non-asymptotic analysis of utility loss and complexity (Theorem 2), demonstrating that the utility loss is optimal and that the iteration complexity improves over first-order methods.
@@
    We evaluate our method on the benchmark dataset. Experiment results show that our algorithm improves the model accuracy, and at the same time saves communication costs compared to Fed-SGD under the same DP guarantee.
@@ 2.1 Basic Setup
The optimization problem (1) satisfies the following conditions:
  1. (i)
    The parameter set $\mathcal{X}$ has finite diameter $D$.
  2. (ii)
    The loss function $l(\cdot,\zeta)$ is $L_{0}$-Lipschitz, $L_{1}$-smooth, and has an $L_{2}$-Lipschitz Hessian for any $\zeta$ over $\mathcal{X}$.
  3. (iii)
    The loss function $l(\cdot,\zeta)$ is $\mu$-strongly convex for any $\zeta$ over $\mathcal{X}$.
From Assumption 1, we infer that also $f_{i}(\cdot)$ and $f(\cdot)$ are $\mu$-strongly convex, $L_{0}$-Lipschitz, $L_{1}$-smooth, and have $L_{2}$-Lipschitz Hessian over $\mathcal{X}$.
@@ 2.2 Newton Methods with Cubic Regularization
Newton methods (Boyd & Vandenberghe (2004)) iteratively minimize a quadratic approximation of the function $f(\cdot)$ as
@@
The cubic regularized Newton method, initially introduced by Nesterov & Polyak (2006), incorporates a second-order Taylor expansion with a cubic regularization term. In particular, the update is
@@
where $L_{2}$ is the Lipschitz Hessian constant in Assumption 2. The cubic upper bound of $f(x_{t})$ in (3) serves as a universal upper bound regardless of the specific characteristics of the objective function. However, the function to minimize in each step of (3) does not have a closed-form solution and it is limited to a centralized single node setting, which our algorithm addresses in a federated setting as discussed in Section 3. Cubic regularization ensures globally convergent second-order optimization with adaptive step control, avoiding the instability and exact Hessian inversion requirements of Newton’s methods while maintaining efficiency in non-convex or distributed settings.
@@ 2.3 Threat Model and DP
Local datasets contain sensitive user information. If problem (1) is addressed in an insecure environment, information leakage could jeopardize both personal and property privacy. This paper considers the following adversary model:
@@
    an honest-but-curious central server that follows the protocol but may attempt to infer private client information from the received messages.
@@
    colluding clients or clients collaborating with the central server to deduce private information about other legal clients.
@@
    an outside eavesdropper who intercepts all transmitted messages without actively destroying communication.
Our adversary model is much stronger than some works that require a trusted third party (Hao et al. (2019); Chen et al. (2024)).
DP is a widely used concept for quantifying privacy risk. It ensures that the presence or absence of any individual in a dataset cannot be inferred from the output of a randomized algorithm $\mathcal{A}$ (Dwork (2006)). Below, we present the formal definition of DP within the context of FL.
@@
Definition 2 states that the output distributions of neighboring datasets exhibit small variation. The factor $\varepsilon$ in (4) represents the upper bound of privacy loss by algorithm $\mathcal{A}$, and $\delta$ denotes the probability of breaking this bound. Therefore, a smaller $\varepsilon$ corresponds to a stronger privacy guarantee. Both Laplace and Gaussian noise can achieve DP. However, Gaussian noise, with its more concentrated distribution and superior composition properties, offers a better balance between privacy and accuracy. Therefore, we focus on the Gaussian mechanism in this work.
@@
(Gaussian Mechanism (Balle & Wang (2018))) A Gaussian mechanism $\mathcal{G}$ for a vector-valued computation $r:\zeta\to\mathbb{R}^{d}$ is obtained by computing the function $r$ on the input data $\zeta_{i}\in\zeta$ and then adding random Gaussian noise perturbation $\nu\sim\mathcal{N}(0,\sigma^{2}I_{d})$ to the output, i.e,
@@ 3 Main Algorithm
In general, there are two approaches for integrating sparsification and privacy in FL: (1) perturb first, then sparsify, and (2) sparsify first, then perturb. The first approach is direct and adaptable since sparsification preserves DP and integrates smoothly with all current privacy mechanisms. However, in the second approach, perturbation may compromise the communication savings achieved through sparsification. Furthermore, empirical observations suggest that the first approach outperforms the second one in some scenarios (Ding et al. (2021)). Therefore, we adopt the first approach in this study.
@@
Following the update of the local model parameter, each client uploads its model update $x_{i,t+1}-x_{t}$ to the server. To address the communication challenges in uplink transmissions, the random-$k$ sparsifier is employed to reduce the size of the transmitted message by a factor of $k/d$ (Li & Richtárik (2021)):
@@
As pointed out by Lacoste-Julien et al. (2012), the output of Algorithm 2 can be computed online. Specifically, setting $z_{0}=\theta_{0}$, and recursively defining $z_{s}=\rho_{s}\theta_{s}+(1-\rho_{s})z_{s-1}$ for $s\geq 1$, with $\rho_{s}=\frac{2}{s+1}$. It is a straightforward calculation to check that $z_{\tau}=\sum_{s=0}^{\tau-1}\frac{2s}{\tau(\tau+1)}\theta_{s}$.
@@
Solving (3) directly requires significant computational resources. To improve efficiency, we use parallel cooperative solving across multiple clients. Since (5) lacks a closed-form solution, we propose a local training approach for its resolution. Privacy is preserved through noise perturbation during local training, while sparsification in uplink transmission reduces communication costs and further enhances privacy protection. Unlike existing methods that transmit compressed Hessian matrices Safaryan et al. (2021); Zhang et al. (2024), we transmit compressed vectors to further minimize communication overhead.